The sum of two numbers is 1974.53. Their difference is 227.23. what are the numbers

Answer by Adrish(44)

SOLUTIONS: Let the number be x and y
Given: x+y=1974.53..............(1)
and x-y=227.23...................(2)
Adding equation (1) and equation(2) we get,
2x=2201.76
x= 2201.76/2
x=1100.88
Substitute the value of x in equation(1) we get
1100.88+y=1974.53
y=1974.53-1100.88
y=873.65
Therefore the numbers are 1100.88 and 873.65